### v2

@@ Abstract
In the field of chemical structure recognition, the task of converting molecular images into graph structures and SMILES string stands as a significant challenge, primarily due to the varied drawing styles and conventions prevalent in chemical literature. To bridge this gap, we proposed MolNexTR, a novel image-to-graph deep learning model that collaborates to fuse the strengths of ConvNext, a powerful Convolutional Neural Network variant, and Vision-TRansformer. This integration facilitates a more nuanced extraction of both local and global features from molecular images. MolNexTR can predict atoms and bonds simultaneously and understand their layout rules. It also excels at flexibly integrating symbolic chemistry principles to discern chirality and decipher abbreviated structures. We further incorporate a series of advanced algorithms, including improved data augmentation module, image contamination module, and a post-processing module to get the final SMILES output. These modules synergistically enhance the model’s robustness against the diverse styles of molecular imagery found in real literature. In our test sets, MolNexTR has demonstrated superior performance, achieving an accuracy rate of 81-97%, marking a significant advancement in the domain of molecular structure recognition.
@@ 1 Introduction
In recent years, with the widespread development of deep neural networks, the performance of optical recognition tasks has significantly improved. However, the recognition of graphical or weakly structured information, such as molecular structure images, remains a challenging problem. In chemical literatures, molecules are usually represented in the form of 2D images. First, the drawing styles of molecules (atomic label fonts, bond drawing styles, etc.) are very diverse and not fully standardized among publishers. Fig.1 shows three different drawing styles of the same molecule. Second, molecules are often drawn as Markush structures. There are no general guidelines for the Markush structure, which leads to much variation in the Markush representation. Thirdly, there are instances where authors of chemical papers employ an artistic flair in representing chemical structures. At times, even seasoned chemists may find themselves at odds grappling with some of these extreme cases. For the reasons mentioned above, the existence of a wide variety of styles in molecular images makes the task of translating these images into machine-understandable molecular structures complex and challenging.
@@
Previous work has demonstrated that it is possible to train CNN(Convolutional Neural Network)-based or ViT(Vision-Transformer)-based models [1, 2, 3, 4, 5, 6, 7, 8] to perform well in some specific styles. However, their robustness and generalization are not guaranteed across so many drawing styles. Moreover, it is impossible to guarantee that the training data contains all possible styles and patterns. Furthermore, ViT-based methods have limitations in translation invariance and local feature representation, and CNN-based methods have limitations in global feature representation. Both local and global features are crucial components for molecular structure recognition, since both the information of the atoms themselves and the information between the atoms are important. In the medical imaging field, some advanced CNN-ViT hybrid architectures have been proposed to address this problem [9, 10, 11, 12, 13] . However, such methods have not yet emerged in the field of molecular structure recognition.
In this work, We aim to enhance the robustness and generalization of the molecular structure recognition model by enhancing the feature extraction ability of the model and the augmentation strategy, to deal with any molecular images that may appear in the real literature. We propose MolNexTR, which is defined as a graph generation model. The model follows the encoder-decoder architecture, takes three-channel molecular images as input, outputs molecular graph structure prediction, and can be easily converted to SMILES. The detailed steps are as follows. First, in the data pre-processing stage, we propose an image contamination augmentation strategy to simulate the interference molecular images (containing other words or other molecular fragments) from real literature. We also propose a series of data augmentation strategies to generate images with different drawing styles during training. Second, in the encoder stage, We propose a novel combined CNN and ViT molecular image encoder network that is aware of both local atom information representation and long-range interatomic feature dependencies in the learning process. Third, in the decoder stage, we combine the Pix2seq and Relationformer architecture [14, 15], using an autoregressive decoder to predict atoms and their coordinates as a sequence and predict the chemical bonds between the atoms, composing the 2D molecular graph. Finally, we include chemical knowledge as symbolic constraints to the model, such as determining the chirality of atoms from the predicted pattern. So that it can accurately recognize complex chemical molecules. An algorithm is used to resolve abbreviated functional clusters commonly found in molecular images. We also designed an error correction algorithm to improve the accuracy of the model for more complex abbreviated functional groups.
@@
Experiments show that MolNexTR outperforms prior image-to-SMILES or image-to-graph-based models on both in-domain and out-of-domain images and achieves strong recognition accuracy (81-97%) on five public benchmarks. Meanwhile, on a previously constructed new benchmark containing molecular images from journal publications, MolNexTR significantly outperforms the existing SOTA, which illustrates the excellent generalization of MolNexTR. In addition, MolNexTR is more robust than SOTA to other disturbances such as low-quality images and input perturbations. The main contributions of this paper are:
@@
  • We propose a transformer-based decoder network with two predictive tasks that enables the model to not only identify the components of molecules but also understand the complex layout rules between them.
@@
  • Our proposed MolNexTR uniquely integrates deep learning model-based techniques with chemical rule-based methods in post-processing part. This enables effective processing of molecular images and the enforcement of chemical constraints during inference.
  • Our proposed MolNexTR demonstrates exceptional performance on multiple challenging datasets including Indigo, ChemDraw, RDKit, CLEF, UOB, USPTO, Staker and ACS.
@@ 2 Related work
Molecular structure identification is a key task in the fields of cheminformatics and computational chemistry with origins dating back to the last century [16]. It is also know as Optical Chemical Structure Recognition(OCSR). Initially, scholars relied on conventional image-processing methodologies and rule-based systems to discern and analyze chemical structures. These rudimentary systems [17, 18, 19, 20, 21, 22, 23, 24, 25, 26] hinged on image binarization, line smoothing, and vectorization techniques, alongside Optical Character Recognition (OCR) to identify atoms and bonds within molecular diagrams. Nonetheless, these frameworks typically necessitated manual rule crafting and heuristic approaches by chemists to tackle diverse scenarios, constraining their widespread applicability and accuracy to some extent. As time progressed, the open-source community contributed several tools [27, 17], addressing newfound challenges in molecular recognition, such as bridge bond identification. Though these systems demonstrated commendable performance on patent imagery, their accuracy dwindled when dealing with the multifaceted imagery of journal articles.
With the rapid evolution of deep learning and neural network technologies in recent years, researchers have embarked on exploring novel methodologies for molecular structure recognition. Notably, by employing Convolutional Neural Networks (CNN) and Recurrent Neural Networks (RNN), researchers devised a novel image-to-SMILES string generation model [1], facilitating automated recognition of molecular structures. After that, a plethora of neural network-based model architectures, such as the Inception network [2], ResNet [8], Transformer [4], Swin Transformer [5, 3] and pretrained-decoder [7], have been proposed to augment the recognition accuracy and robustness of these systems.
@@
To mitigate these issues, novel systems like ChemGrapher [28] and MolMiner [29] have adopted distinctive approaches. They train separate modules to detect atoms, bonds, and texts, followed by heuristic-based graph construction, thus allowing the integration of chemical constraints during the construction phase.  [3, 6] have adopted end-to-end model to generate molecular graphs. This approach eschews reliance on heuristic methods for connecting local predictions, simplifying the model architecture, and paving new avenues for the future evolution of molecular structure recognition technology.
@@ 3.1 Problem Definition
Molecular structure recognition aims to transfer images of individual molecules into their corresponding molecular structures with information of atoms and bonds. In our work, We define this problem as an image-to-sequence formulation. Given an image $I$ of a single molecular, we transfer it into graph sequence $S^{G}=(S^{M},S^{N})$ that represents molecular structures, where $S^{M}=[M_{1},M_{2},...,M_{n}]$ is the atoms sequence and $S^{N}=[N_{1},N_{2},...,N_{n}]$ is the bonds sequence. Specifically, $M_{n}=(l_{n},x_{n},y_{n})$ is the definition of each atom, $l_{n}$ is the corresponding SMILES string of the atom and $(x_{n},y_{n})$ is the 2D coordinates of the atom in the image. $N_{n}=[m_{n,1}T,m_{n,2}T,...,m_{n,n}T]$ is the set of bonds of each atom, $m_{n,1}T$ means the $n$-th atom is connected to the $1st$ atom with bond $T$, where $T$ is the set of bonds type that represents a different type of bonds (eg., single, double, solid wedge, dashed wedge). Finally, it can be easily converted into molecular graphs and standard data.
@@ 3.2.3 Molecular Augmentation
Figure 2: The illustration of The Molecular Augmentation. The Molecular Augmentation consists of four main actions.1) replace functional group, 2) add complex abbreviations, 3) add C bond, 4) add R-group.
In most chemistry literature, authors map molecules with various functional groups abbreviations and R-group substituents. However, when rendering images using RDKit or Indigo, functional groups or R-groups are never included. To generate molecules with such abbreviations and substituents, we created a list of more than 100 common functional groups with their corresponding abbreviations. Augmenting algorithms randomly replace functional groups with abbreviations in molecules to generate augmented datasets. Specifically, in the process of molecular augmentation, if a functional group exists in a molecule, we randomly replace this functional group with the corresponding abbreviation in the list according to a given probability. The original structure branch is removed from the molecular graph. To provide a variety of R-groups, we also have a list of R-group labels and randomly add them to the molecule. We also randomly added noise keys to simulate possible interference in real images. Furthermore, another possible style of abbreviations is chain abbreviations such as CH3CH2NH2, which are longer and more complex than the abbreviations in the list and are impossible to list completely, but converting them to SMILES is easy as long as the right characters are identified. Therefore, for the model to generalize better to those not covered in the list. We also have a collection of chained abbreviation components (eg. CH3, CH2, NH2, OH) that are randomly combined to form a complex abbreviation. We believe that this method can improve the model capability of optical character recognition (OCR) so that the model can recognize such unseen abbreviations and will not ignore abbreviations whose length and complexity are far beyond the known abbreviations in the list. The process of our molecular augmentation is illustrated in Fig. 2.
@@ 3.2.4 Image Contamination Algorithm
Figure 3: The illustration of Image Contamination Algorithm. The Contamination Algorithm consists of six main actions which are 1) add atom noise, 2)add bond noise, 3) add incomplete structural noise, 4) add line noise, 5) add incomplete atom noise, and 6) add arrow noise.
@@ 3.3 Network Overview
Figure 4: Overview of our MolNexTR model. The molecular image is first encoded with the dual-stream encoder. Then a transformer-based decoder is applied for sequential atom and bond prediction. Finally, the post-processing modules ensures that the various molecular structures can be accurately converted to SMILES, SMART or MOLfile formats..
The Overview of our MolNexTR is illustrated in Fig. 4. It follows the encoder-decoder architecture and consists of three main components: 1) For the encoder network, we propose a dual-stream encoder (ref. Sec. 3.4), which concurrently combines a CNN and a vision transformer network to capture local feature dependencies and long-range feature dependencies, respectively. Such a design can avoid the problems of incomplete atom information (lack of local features) and incomplete relationships between atoms(lack of global features) as much as possible.; 2) After the input image is encoded into a feature representation, we use a transformer-based decoder with a two-stage prediction(atom prediction and bond prediction) as the structure decoder network, so that our decoder can predict each atom and bond in the molecule, as well as their geometric arrangement. 3) Afterward, we employ a post-processing module to ensure that the model can accurately and comprehensively reconstruct structures of various molecular styles. Finally, these structures can be easily converted into a SMILES string [32], SMARTS or MOLfile [33].
@@ 3.4.2 The ViTs Stream
In addition to short-range feature dependencies, long-range feature dependencies between image patches also matter for molecular recognition. To this end, the long-range feature dependencies in different feature scales are captured by the vision transformer (ViT) [36], which is composed of multiple parallel transformer blocks that receive feature patches under different scales as input. All transformer blocks share a similar structure, consisting of patch embedding layers and transformer encoding layers. As illustrated in the bottom of Fig. 4(a), it shows the implementation process of ViT on the input feature map $\text{F}_{1}^{c}\in\mathbb{R}^{\frac{H}{4}\times\frac{W}{4}\times C}$. Firstly, the input feature map is divided into $\frac{HW}{16p^{2}}$ patches with size $p\times p$, and each patch is flattened into a vector $\mathbf{v}_{n}\in\mathbb{R}^{p^{2}\times C}$. In this work, four parallel transformer blocks are used, which receive feature patches of size $p=4,8,16,32$. Then, a linear projection layer is applied to each patch vector to obtain the patch embedding $\text{e}_{n}\in\mathbb{R}^{C}$. The patch embeddings, together with position embeddings, are then fed into the transformer encoding layers to obtain the output. Each encoding layer comprises a lightweight multi-head self-attention (MHSA) layer and a feed-forward network (FFN). The MHSA layer receives a truncated query $Q$, key $K$, and value $V$ as input and computes the attention score $\text{A}\in\mathbb{R}^{N\times N}$ as follows:
@@ 3.6.1 Stereochemical discrimination module
The stereochemistry of organic molecules exerts a significant impact on their structure and properties. However, recognizing stereochemistry poses challenges for neural models. In SMILES notation, chirality is indicated as an atomic property, with ”@” or ”@@” following an atom label to describe the relative spatial arrangement of connected bonds, following their order of listing. There’s no direct mapping between image patterns and the presence of ”@” or ”@@”. Furthermore, if there is a functional group around the chiral center, it becomes more difficult to discriminate.

### v3

@@ Abstract
In the field of chemical structure recognition, the task of converting molecular images into machine-readable data formats such as SMILES string stands as a significant challenge, primarily due to the varied drawing styles and conventions prevalent in chemical literature. To bridge this gap, we proposed MolNexTR, a novel image-to-graph deep learning model that collaborates to fuse the strengths of ConvNext, a powerful Convolutional Neural Network variant, and Vision-TRansformer. This integration facilitates a more detailed extraction of both local and global features from molecular images. MolNexTR can predict atoms and bonds simultaneously and understand their layout rules. It also excels at flexibly integrating symbolic chemistry principles to discern chirality and decipher abbreviated structures. We further incorporate a series of advanced algorithms, including an improved data augmentation module, an image contamination module, and a post-processing module for getting the final SMILES output. These modules cooperate to enhance the model’s robustness to diverse styles of molecular images found in real literature. In our test sets, MolNexTR has demonstrated superior performance, achieving an accuracy rate of 81-97%, marking a significant advancement in the domain of molecular structure recognition.
@@ 1 Introduction
In recent years, with the widespread development of deep neural networks, the performance of optical recognition tasks has significantly improved. However, the recognition of graphical or weakly structured information, such as molecular structure images, remains a challenging problem. In chemical literatures, molecules are usually represented in the form of 2D images. First, the drawing styles of molecules (atomic label fonts, bond drawing styles, etc.) are very diverse and not fully standardized among publishers. Fig. 1 shows three different drawing styles of the same molecule. Second, molecules are often drawn as Markush structures. There are no general guidelines for the Markush structure, which leads to much variation in the Markush representation. Third, there are instances where authors of chemical papers employ an artistic flair in representing chemical structures. Even experienced chemists can sometimes struggle to understand these more unusual representations. For the reasons mentioned above, the existence of a wide variety of styles in molecular images makes the task of translating these images into machine-understandable molecular structures complex and challenging.
@@
Previous work has demonstrated that it is possible to train convolutional neural network (CNN)-based or vision-transformer (ViT)-based models [1, 2, 3, 4, 5, 6, 7, 8] to perform well in some specific styles. However, their robustness and generalization are not guaranteed across so many drawing styles. Moreover, it is impossible to guarantee that the training data contains all possible styles and patterns. Furthermore, ViT-based methods have limitations in translation invariance and local feature representation, and CNN-based methods have limitations in global feature representation. Both local and global features are crucial components for molecular structure recognition, since both the information of the atoms themselves and the information between the atoms are essential. In the medical imaging field, some advanced CNN-ViT hybrid architectures have been proposed to address this problem [9, 10, 11, 12, 13]. However, such methods have not yet emerged in the field of molecular structure recognition.
In this work, we aim to enhance the robustness and generalization of the molecular structure recognition model by enhancing its feature extraction ability and augmentation strategies, which can deal with any molecular images that may appear in the real literature. We propose MolNexTR (Molecular convNext-TRansformer), which is defined as a graph generation model. The model follows the encoder-decoder architecture, takes three-channel molecular images as input, outputs molecular graph structure prediction, and can be easily converted to machine-readable data formats such as simplified molecular-input line-entry system (SMILES) [14]. The detailed steps are as follows. First, in the data pre-processing stage, we propose an image contamination augmentation strategy to simulate the interference molecular images (containing other words or other molecular fragments) from real literature. We also propose a series of data augmentation strategies to generate images with different drawing styles during training. Second, in the encoder stage, we propose a novel combined CNN and ViT molecular image encoder network that is aware of both local atom information representation and long-range interatomic feature dependencies in the learning process. Third, in the decoder stage, we combine the Pix2seq and Relationformer architecture [15, 16], using an autoregressive decoder to predict atoms and their coordinates as a sequence and predict the chemical bonds between the atoms, composing the 2D molecular graph. Finally, we include chemical knowledge as symbolic constraints to the model, such as determining the chirality of atoms from the predicted pattern, so that it can accurately recognize complex chemical molecules. An algorithm is used to resolve abbreviated functional groups commonly found in molecular images. We also designed an error correction algorithm to improve the accuracy of the model for more complex abbreviated functional groups.
@@
Experiments show that MolNexTR outperforms prior image-to-SMILES or image-to-graph-based models on both in-domain and out-of-domain images and achieves strong recognition accuracy (81-97%) on five public benchmarks. Meanwhile, on a previously constructed new benchmark containing molecular images from journal publications, MolNexTR significantly outperforms the existing methods, which illustrates the excellent generalization of MolNexTR. In addition, MolNexTR is more robust than existing methods in handling other disturbances such as low-quality images and input perturbations. The main contributions of this paper are:
@@
  • We propose a transformer-based decoder network with two predictive tasks that enable the model to not only identify the components of molecules but also understand the complex layout rules between them.
@@
  • Our proposed MolNexTR uniquely integrates deep learning model-based techniques with chemical rule-based methods in the post-processing part. This enables the effective processing of molecular images and the enforcement of chemical constraints during inference.
  • Our proposed MolNexTR demonstrates exceptional performance on multiple challenging datasets including Indigo, ChemDraw, RDKit, CLEF, UOB, JPO, USPTO, Staker and ACS.
@@ 2 Related work
Molecular structure identification is a key task in the fields of cheminformatics and computational chemistry with origins dating back to the last century [17]. It is also known as optical chemical structure recognition (OCSR). Initially, scholars relied on conventional image-processing methodologies and rule-based systems to discern and analyze chemical structures. These basic systems [18, 19, 20, 21, 22, 23, 24, 25, 26, 27] depended on image binarization, line smoothing, vectorization techniques, and optical character recognition (OCR) to identify atoms and bonds within molecular images. Nonetheless, these frameworks typically needed manual rules and heuristic approaches created by chemists to tackle diverse scenarios, constraining their widespread applicability and accuracy to some extent. As time progressed, the open-source community contributed several tools [28, 18], addressing newfound challenges in molecular recognition, such as bridge bond identification. Though these systems demonstrated commendable performance on patent images, their accuracy decreased when dealing with the multifaceted images of journal articles.
With the rapid evolution of deep learning and neural network technologies in recent years, researchers have embarked on exploring novel methodologies for molecular structure recognition. Notably, by employing a convolutional neural network and a recurrent neural network (RNN), researchers devised a novel image-to-SMILES string generation model [1], facilitating automated recognition of molecular structures. After that, a plethora of neural network-based model architectures, such as the Inception network [2], ResNet [8], Transformer [4], Swin Transformer [5, 3] and pretrained-decoder [7], have been proposed to augment the recognition accuracy and robustness of these systems.
@@
To mitigate these issues, novel systems like ChemGrapher [29] and MolMiner [30] have adopted distinctive approaches. They train separate modules to detect atoms, bonds, and texts, followed by heuristic-based graph construction, thus allowing the integration of chemical constraints during the construction phase. Qian et al. and Yoo et al. have adopted an end-to-end model to generate molecular graphs [3, 6]. This approach eschews reliance on heuristic methods for connecting local predictions, simplifying the model architecture, and paving new avenues for the future evolution of molecular structure recognition technology.
@@ 3.1 Problem Definition
Molecular structure recognition aims to transfer images of individual molecules into their corresponding molecular structures with information of atoms and bonds. In our work, we define this problem as a graph sequence formulation. Given an image $I$ of a single molecular, we transfer it into graph sequence $S^{G}=(S^{M},S^{N})$ that represents molecular structures, where $S^{M}=[M_{1},M_{2},...,M_{n}]$ is the atoms sequence and $S^{N}=[N_{1},N_{2},...,N_{n}]$ is the bond sets sequence. Specifically, $M_{n}=(l_{n},x_{n},y_{n})$ is the definition of each atom, $l_{n}$ is the corresponding SMILES string of the atom and $(x_{n},y_{n})$ is the 2D coordinates of the atom in the image. $N_{n}=[m_{n,1}T,m_{n,2}T,...,m_{n,n}T]$ is the set of bonds of each atom, $m_{n,1}T$ means the $n$-th atom is connected to the $1st$ atom with bond $T$, where $T$ is the set of bond types that represents a different type of bonds (e.g., single, double, solid wedge, dashed wedge, or none). Finally, it can be easily converted into molecular graphs and standard data formats.
@@ 3.2.3 Molecular Augmentation
Figure 2: The illustration of the molecular augmentation. The Molecular Augmentation consists of four main actions: 1) replace functional group, 2) add complex abbreviations, 3) add C bond, 4) add R-group.
In most chemistry literature, authors map molecules with various functional groups abbreviations and R-group substituents. However, when rendering images using RDKit or Indigo, functional groups or R-groups are never included. To generate molecules with such abbreviations and substituents, we created a list of more than 100 common functional groups with their corresponding abbreviations. Augmenting algorithms randomly replace functional groups with abbreviations in molecules to generate augmented datasets. Specifically, in the process of molecular augmentation, if a functional group exists in a molecule, we randomly replace this functional group with the corresponding abbreviation in the list according to a given probability. The original structure branch is removed from the molecular graph. To provide a variety of R-groups, we also have a list of R-group labels and randomly add them to the molecule. We also randomly added noise keys to simulate possible interference in real images. Furthermore, another possible style of abbreviations is chain abbreviations such as CH3CH2NH2, which are longer and more complex than the abbreviations in the list and are impossible to list completely, but converting them to SMILES is easy as long as the right characters are identified. Therefore, for the model to generalize better to those not covered in the list, we also have a collection of chained abbreviation components (e.g., CH3, CH2, NH2, OH) that are randomly combined to form a complex abbreviation. We believe that this method can improve the model capability of OCR so that the model can recognize such unseen abbreviations and will not ignore abbreviations whose length and complexity are far beyond the known abbreviations in the list. The process of our molecular augmentation is illustrated in Fig. 2.
@@ 3.2.4 Image Contamination Algorithm
Figure 3: The illustration of the image contamination algorithm. The Contamination Algorithm consists of six main actions which are 1) add atom noise, 2) add bond noise, 3) add incomplete structural noise, 4) add line noise, 5) add incomplete atom noise, and 6) add arrow noise.
@@ 3.3 Network Overview
Figure 4: Overview of our MolNexTR model. The molecular image is first encoded with the dual-stream encoder. Then a transformer-based decoder is applied for sequential atom and bond prediction. Finally, the post-processing modules ensure that the various molecular structures can be accurately converted to SMILES, SMART or MOLfile formats.
The Overview of our MolNexTR is illustrated in Fig. 4. It follows the encoder-decoder architecture and consists of three main components: 1) For the encoder network, we propose a dual-stream encoder (ref. Sec. 3.4), which concurrently combines a CNN and a vision transformer network to capture local feature dependencies and long-range feature dependencies, respectively. Such a design can avoid the problems of incomplete atom information (lack of local features) and incomplete relationships between atoms (lack of global features) as much as possible.; 2) After the input image is encoded into a feature representation, we use a transformer-based decoder with a two-stage prediction (atom prediction and bond prediction) as the structure decoder network, so that our decoder can predict each atom and bond in the molecule, as well as their geometric arrangement. 3) Afterward, we employ a post-processing module to ensure that the model can accurately and comprehensively reconstruct structures of various molecular styles. Finally, these structures can be easily converted into a SMILES string, a smiles arbitrary target specification (SMARTS) string, or a MOLfile [33].
@@ 3.4.2 The ViTs Stream
In addition to short-range feature dependencies, long-range feature dependencies between image patches also matter for molecular recognition. To this end, the long-range feature dependencies in different feature scales are captured by the ViTs stream [36], which is composed of multiple parallel transformer blocks that receive feature patches under different scales as input. All transformer blocks share a similar structure, consisting of patch embedding layers and transformer encoding layers. As illustrated in the bottom of Fig. 4(a), it shows the implementation process of ViT on the input feature map $\text{F}_{1}^{c}\in\mathbb{R}^{\frac{H}{4}\times\frac{W}{4}\times C}$. Firstly, the input feature map is divided into $\frac{HW}{16p^{2}}$ patches with size $p\times p$, and each patch is flattened into a vector $\mathbf{v}_{n}\in\mathbb{R}^{p^{2}\times C}$. In this work, four parallel transformer blocks are used, which receive feature patches of size $p=4,8,16,32$. Then, a linear projection layer is applied to each patch vector to obtain the patch embedding $\text{e}_{n}\in\mathbb{R}^{C}$. The patch embeddings, together with position embeddings, are then fed into the transformer encoding layers to obtain the output. Each encoding layer comprises a lightweight multi-head self-attention (MHSA) layer and a feed-forward network (FFN). The MHSA layer receives a truncated query $Q$, key $K$, and value $V$ as input and computes the attention score $\text{A}\in\mathbb{R}^{N\times N}$ as follows:
@@ 3.6.1 Stereochemical discrimination module
The stereochemistry of organic molecules exerts a significant impact on their structure and properties. However, recognizing stereochemistry poses challenges for neural models. In SMILES notation, chirality is indicated as an atomic property, with "@" or "@@" following an atom label to describe the relative spatial arrangement of connected bonds, following their order of listing. There’s no direct mapping between image patterns and the presence of "@" or "@@". Furthermore, it becomes even more challenging to discriminate when there are complex functional groups connected to the chiral center.